OpenStudy (saifoo.khan):

Determine the zeros of f(x) = x3 – 12x2 + 28x – 9.

OpenStudy (anonymous):

YOU need an answer !? I'm surprised lol ...

OpenStudy (saifoo.khan):

LOL.

OpenStudy (saifoo.khan):

Method as well.

OpenStudy (chaise):

When you mean zeros do you mean the points of intersection of the graph?

OpenStudy (saifoo.khan):

idk, that''s all wht the question says. i guess that's same.

OpenStudy (anonymous):

There is a method

OpenStudy (chaise):

y = x^3 – 12x^2 + 28x – 9. When the graph cuts the x axis y is equal to zero. 0 = x^3 – 12x^2 + 28x – 9. Solve it using whatever method you like. x = 9 x = 0.381966 x = 2.61803 I hope what I think your asking is what I'm answering. What is your chapter on?

OpenStudy (anonymous):

There's even a formula! (similar to the general solution for quadratic equations), but it's much much much longer :-P

OpenStudy (anonymous):

I'll post the method (which I forgot ages ago)

OpenStudy (saifoo.khan):

now @chayC. np man!

OpenStudy (saifoo.khan):

how*

OpenStudy (saifoo.khan):

?

OpenStudy (anonymous):

try completing the cube,

OpenStudy (saifoo.khan):

hmm

OpenStudy (anonymous):

f(x) = x3 – 12x2 + 28x – 9. I don't think you can easily complete the cube, so we will use the long method

OpenStudy (anonymous):

substitute x for (t - (b/(3a))), which is t + 4 (t+4)^2 - 12(t+4)^2 + 28(t+4) - 9 = 0

OpenStudy (anonymous):

oops it's: (t+4)^3 - 12(t+4)^2 + 28(t+4) - 9 = 0

OpenStudy (anonymous):

expand to get t^3 -20t -25 = 0

OpenStudy (saifoo.khan):

my czn showed me this method, it was long as hell. thanks for your help. :0

OpenStudy (anonymous):

it's easier than using the general solution :-P

OpenStudy (anonymous):

after all that.. I give up

OpenStudy (anonymous):

I think the best way would be to guess the first rational root

OpenStudy (saifoo.khan):

lol, nvm.. thanks for your time.. :)

OpenStudy (anonymous):

the way to do it is to take the last term, which has factors 2, 3, 9

OpenStudy (anonymous):

oh and 1

OpenStudy (anonymous):

signed and unsigned

OpenStudy (saifoo.khan):

how about solving it by taking commons?

OpenStudy (anonymous):

plug in those values, and see which one gives you zero for the function

OpenStudy (anonymous):

I got 9, so we can factor out (x-9) use long division to find out what the rest of the equation looks like after you factor out x-9

OpenStudy (saifoo.khan):

that's right.. but how to get x-9 with taking commons?

OpenStudy (anonymous):

(x-9)(x^2-3x+1) now it's easy since you can use the general formula for quadratic roots

OpenStudy (anonymous):

taking commons..... hmm

OpenStudy (saifoo.khan):

i can solve the right part with using quadratic formula. but how you got x-9?

OpenStudy (anonymous):

trial and error :-D

OpenStudy (saifoo.khan):

Cool. Thanks.

OpenStudy (anonymous):

best to memorize that general method (or the general formula for cubic roots if you're guinness world record genius, but you will be able to derive it if you know the procedure)

OpenStudy (saifoo.khan):

Okay, thanks!! :) i have it in written form, but havn't memorized it.

OpenStudy (anonymous):

There is a method that involves complex numbers. I haven't tried it but it might be easier. http://en.wikipedia.org/wiki/Cubic_function#Lagrange.27s_method

OpenStudy (anonymous):

let \[\zeta =-\frac{1}{2}+\frac{\sqrt{3}}{2}i\]so that zeta^2 + zeta + 1 = 0

OpenStudy (anonymous):

now, set \[s_0 = x_0+x_1+x_2\] \[s_1 = x_0 + \zeta x_1+\zeta^2x_2\] \[s_2=x_0 +\zeta^2x_1+\zeta x_2\]

OpenStudy (anonymous):

where, x_0, x_1, and x_2 are the possible roots of ax^3+bx^2+cx+d=0

OpenStudy (saifoo.khan):

Dang! that's scary. :O

OpenStudy (saifoo.khan):

lol. that's Lagrange's ,ethod.

OpenStudy (saifoo.khan):

method*

OpenStudy (anonymous):

now you can do the reverse Fourier Transform to get the real roots \[x_0 = \frac{1}{3}(s_0+s_1+s_2)\] \[x_1 = \frac{1}{3}(s_0+\zeta^2s_1+\zeta s_2)\] \[x_2 = \frac{1}{3}(s_0 +\zeta s_1+\zeta^2s_2)\]

OpenStudy (anonymous):

finding s_0 is straightforward \[s_0=-\frac{b}{a}=12\] finding s_1 and s_2 is more involved

OpenStudy (anonymous):

trololololol. he's asking about the zeros of the function, either descartes' or rational theorem. he's prolly in the fundamental theorem of algebra.